OpenStudy (pooh):

Train A and Train B are traveling in the same direction going parallel. Train a is traveling at 40 miles an hour and train be is traveling at 60 miles per hour. Train A passes the station at 7:20 am and B passes at 7:34, what time will train b catch up

OpenStudy (anonymous):

Trains A= 40t+14/60 * 40 Trains B=60t solve for t 40t+28/3=60t 28/3=20t 28=60t t=7/15

OpenStudy (anonymous):

7/15 hours is 28 minutes add that to 7:34

OpenStudy (pooh):

so is that 7 hours and 15 minutes later?? I am totally confused I feel that its not giving enough information

OpenStudy (pooh):

OKAY SO THIS MEANS THAT B WILL CATCH UP AT 8:02

OpenStudy (anonymous):

\[\frac{7}{15}\] hours is 28 minutes example \[\frac{1}{2}\] hours is 30 minutes

OpenStudy (anonymous):

8:02 is right

OpenStudy (anonymous):

We multiply by 60 to go from hours to minutes 7/15 *60 =28 minutes